OpenStudy (anonymous):

PLEASE HELP!!!! Find the common solution to the following pair of lines by graphing them on the same axes. 3x + 2y = 6 x - y = 2 The lines intersect at (__, __).

OpenStudy (anonymous):

(2,0)

OpenStudy (anonymous):

x-y=2 is equal to y=x-2 and 3x+2y=6 is equal to y=(-3/2)x+3. Can you graph these?